ZombieSlayer:

Which statement is true regarding the graphed functions? On a coordinate plane, a straight red line with a positive slope, labeled g of x, crosses the x-axis at (negative 6, 0) and the y-axis at (0, 6). A straight blue line with a negative slope, labeled f of x, crosses the x-axis at (negative 0.75, 0) and the y-axis at (0, negative 2). f(4) = g(4) f(4) = g(–2) f(2) = g(–2) f(–2) = g(–2)

ZombieSlayer:

1 attachment
sealake101:

ok so what is not correct?

jimthompson5910:

Where do the two lines cross?

ZombieSlayer:

Im looking for the intersection point?

jimthompson5910:

Yes because the solution to f(x) = g(x) is where the two lines cross. Specifically all we're taking out of that is the x coordinate of that intersection point.

ZombieSlayer:

Ah, thank you!!!

jimthompson5910:

Recall that y = f(x), Also y = g(x) as well So if we had f(x) = g(x), then we're basically asking "when are the y values the same for some x value?"

ZombieSlayer:

Thank You!!

jimthompson5910:

no problem

ZombieSlayer:

So, woulf it be f(4) = g(–2)?

jimthompson5910:

no

jimthompson5910:

f(x) = g(x) must have the same x value you replace every x with the same number

ZombieSlayer:

so not to sound dumb, woud it be D.?

jimthompson5910:

D is correct The two lines intersect at (-2, 4) here x = -2 is plugged into f(x) and g(x) we go from f(x) = g(x) to f(-2) = g(-2)

ZombieSlayer:

Thank you!!!

jimthompson5910:

Notice that f(-2) = 4 and g(-2) = 4 since both are 4, we can equate f(x) and g(x) for this x value

jimthompson5910:

sure thing

ZombieSlayer:

ohhhhh, that makes sense